b.) 5 x1/3/(x1/3 – 1) dx
Answer #1

/ da !! x 1 80, Any = *(4-1)+bm{+-1 +9(+-1) + 9 lt-ig? = (x²21) 4 in (421) + 9133_1] +1 (121 put a = t3 da = 3+2 dt t 3+2 dt